# StyleGuard: Preventing Text-to-Image-Model-based Style Mimicry Attacks by Style Perturbations

### v1

@@ Abstract
Recently, text-to-image diffusion models have been widely used for style mimicry and personalized customization through methods such as DreamBooth and Textual Inversion. This has raised concerns about intellectual property protection and the generation of deceptive content. Recent studies, such as Glaze and Anti-DreamBooth, have proposed using adversarial noise to protect images from these attacks. However, recent purification-based methods, such as DiffPure and Noise Upscaling, have successfully attacked these latest defenses, showing the vulnerabilities of these methods. Moreover, present methods show limited transferability across models, making them less effective against unknown text-to-image models. To address these issues, we propose a novel anti-mimicry method, StyleGuard. We propose a novel style loss that optimizes the style-related features in the latent space to make it deviate from the original image, which improves model-agnostic transferability. Additionally, to enhance the perturbation’s ability to bypass diffusion-based purification, we designed a novel upscale loss that involves ensemble purifiers and upscalers during training. Extensive experiments on the WikiArt and CelebA datasets demonstrate that StyleGuard outperforms existing methods in robustness against various transformations and purifications, effectively countering style mimicry in various models. Moreover, StyleGuard is effective on different style mimicry methods, including DreamBooth and Textual Inversion.
@@ 1 Introduction
Diffusion models have demonstrated remarkable effectiveness across various applications, such as image generation Dhariwal and Nichol (2021); Ho et al. (2020), image editing Kawar et al. (2023); Yang et al. (2023); Zhang et al. (2023), and text-to-image synthesis Xie et al. (2023); Wang et al. (2024c). The emergence of text-to-image generator models, such as Stable Diffusion (SD), has significantly transformed the art industry. These models allow users to create detailed artwork from simple text prompts, a task that once required extensive time and skill from professional artists. However, these technologies have also raised concerns regarding copyrights and ethics. For example, Dreambooth Ruiz et al. (2023) allows anyone to fine-tune a SD model to imitate the artistic style of an artist with just a few paintings and generate high-quality artwork. This seriously damages the intellectual property rights of artists.
@@
To tackle the challenges associated with unauthorized image usage in text-to-image generation, recent perturbation-based approaches have emerged. These methods are designed to subtly modify user images, rendering them "unlearnable" for malicious applications and disrupting the functionality of targeted diffusion models. For example, Anti-Dreambooth Van Le et al. (2023) shows some effectiveness by alternately training diffusion models and executing PGD attacks, but it is fragile against simple data transformations. MetaCloak Liu et al. (2024) builds on Anti-Dreambooth by incorporating simple transformations into the attack pipeline. SimAC Wang et al. (2024a) improves MetaCloak by selecting time steps with the highest gradients to improve the stability of the training.
@@
Although these methods have demonstrated effectiveness against style mimicry, recent purification-based methods pose significant challenges to these protection techniques. For example, DiffPure Nie et al. (2022) and Noise Upscaling Hönig et al. (2025) utilize LDM-based purifiers and upscalers to purify noise and successfully bypass the defenses of these methods, as illustrated in Figure 1. These novel attacks highlight the vulnerabilities of current anti-customization techniques. Moreover, in real-world settings, the fine-tuning model may differ from the model used in the training phase, making cross-model transferability an important issue. In light of these limitations, there is a need for a more efficient method to generate protective noise that is resilient to diffusion-based purifications and has strong cross-model transferability.
@@
To solve these problems, we propose StyleGuard, a more effective method based on style perturbations to protect artists from unauthorized text-to-image diffusion-based style mimicry. To generate transferable perturbations, inspired by style transfer Huang and Belongie (2017a) and adversarial attacks in feature space Xu et al. (2021), we disrupt style-related features in latent space by adding subtle noise to the fine-tuning images that is nearly imperceptible to the human eye. Our method prevents text-to-image models from accurately extracting the style features of fine-tuning images, leading to false style correlations and hindering attackers from mimicking the original image’s style. Moreover, compared to Mist Liang and Wu (2023) that uses $L_{2}$ distance in the feature space, which is prone to causing local disturbances and sensitive to image variations, style perturbations alter global features. As a result, our method is more robust to transformations such as image cropping, rescaling, and Gaussian noise. Additionally, bypassing diffusion-based purifications presents significant challenges, particularly because directly incorporating purifiers or upscalers into the optimization process can lead to memory overflow, as most purifiers and upscalers are based on diffusion models Nie et al. (2022); Lee and Kim (2023); Chen et al. (2024). To address this issue, we propose a novel upscale loss function that maximizes the loss of denoise error on ensemble purifiers and upscalers through a meta-learning approach. On the WikiArt and ClebA-HQ datasets, we show that StyleGuard significantly outperforms existing approaches over various transformations and purification methods and has higher cross-model transferability. Our main contributions are summarized as follows.
  • We propose StyleGuard, the first robust method to protect artists from style mimicry attacks that can further bypass DiffPure and Noise Upscaling defenses.
@@
  • To improve the model-agnostic transferability, we introduce a style loss to generate protective noise. This method aligns the style characteristics of the protected image more closely with those of the target image, making the fine-tuned model to establish an incorrect style connection.
@@
  • We conducted comparisons with state-of-the-art methods on the WikiArt and CelebA datasets. Experiments demonstrate that our method offers enhanced protection, effectively preventing style mimicry for both DreamBooth and Textual Inversion. Moreover, our method is also effective on preventing identity customization.
@@ 2.1 Style Mimicry and Copyright Concerns
Unauthorized style mimicry has become a significant concern in the AI art community, where malicious actors exploit AI models to replicate an artist’s unique style without consent Andersen (2022); Baio (2022); Murphy (2022). Such attacks often begin with a naive approach, where a generic text-to-image model is queried using the name of a well-known artist. More advanced mimicry attacks involve fine-tuning generic text-to-image models on a small collection of an artist’s works, often as few as 20 pieces by models such as DreamBooth Gal et al. (2022a); Ruiz et al. (2023). DreamBooth identifies key stylistic features and associates them with a specific token in the fine-tuned model, enabling highly accurate style replication. Such techniques have led to widespread incidents of unauthorized mimicry Andersen (2022); Baio (2022); Murphy (2022), for example, CivitAI CivitAI (2022) built a large online website where people share their finetuned stable diffusion models. The potential for unauthorized style mimicry threatens the livelihoods of artists, leading to discussions about intellectual property rights in the digital age.
@@ 2.2 Protection Against Style Mimicry and Personalization
To address the unauthorized style mimicry issue, perturbation-based methods have been developed, which add subtle image perturbations to the unprotected images to disrupt generative models. For example, PhotoGuard Salman et al. (2023) aligns protected images’ latent features with black-and-white images. Glaze Shan et al. (2023) minimizes the feature distance between perturbed images and target images while maintaining perceptual similarity. AdvDM Liang et al. (2023) reduces the likelihood of perturbed images under pre-trained diffusion models by disrupting the denoising process. Its enhanced version, Mist Liang and Wu (2023), utilizes black-and-white periodic images as targets and incorporates semantic loss and textual loss to improve protective strength. However, Mist directly minimizes the L2 distance between the original and target latent features, which causes noticeable and unnatural textures that degrade the original image’s quality. However, Mist often results in significant noise, which can greatly degrade the visual quality of the images.
DreamBooth can be used not only for style imitation, but also for generating photos of specific people, which is generally known as personalization attack. Anti-DreamBooth Van Le et al. (2023) proposes a novel scheme to defend the personalization attack that alternately updates the diffusion model and protected images. MetaCloak Liu et al. (2024) builds upon Anti-DreamBooth by incorporating simple transformations, such as Gaussian blur and cropping, into the attack pipeline to improve robustness against these transformations. SimAC Wang et al. (2024a) further extends the work done by Anti-DreamBooth by selecting timesteps with maximum gradients to stabilize the training process. Some other methods Huang et al. (2024); Zhu et al. (2024); Yao et al. (2024); Wang et al. (2024b) protect copyrights by embedding watermarks into images, subtly incorporating the author’s information. However, this kind of approach introduces additional verification processes.
While these methods exhibit robustness to simple transformations, recent work has introduced purification-based methods, including DiffPure Nie et al. (2022) and Noise Upscaling Hönig et al. (2025), which first add noise to images and then employ an LDM as purifier or upscaler to remove noise. These approaches have demonstrated impressive results in removing protective noise, rendering many recent protective methods ineffective. Therefore, there is an urgent need to develop more robust methods to defend against such attacks, particularly in light of the challenges posed by the purification-based attacks.
@@ 3.1 Style Mimicry by Dreambooth
DreamBooth Ruiz et al. (2023) introduces a novel approach for personalizing text-to-image diffusion models by enabling them to generate high-fidelity images of specific subjects based on a few reference images. The method fine-tunes a pre-trained text-to-image model to bind a unique identifier (e.g., "[V]") to the subject, allowing the model to synthesize the subject in diverse contexts while preserving its key visual features. This is achieved through a fine-tuning process: the latent diffusion model (LDM) is fine-tuned using input images paired with text prompts containing the unique identifier and the subject’s class name (e.g., "An [V] painting"), while a class-specific prior preservation loss ensures the model retains its semantic understanding of the broader class (e.g., "A painting"). The DreamBooth loss function is defined as
@@ 4.2 Disturbing Style-related Features and Generating Transferable Perturbations
Previous work has found that the mean and variance of the feature space encapsulate the style information Ulyanov et al. (2016); Xu et al. (2021); Huang and Belongie (2017b); Karras et al. (2019a). We consider the problem defined in Eq. 4.1 as maximizing the style-related feature distance between the reference images and the perturbed images and making it closer to the target images. Therefore, we propose a novel style loss function, which is defined as
@@ 4.3 Why Previous Defenses are Ineffective to Noise Upscaling and DiffPure
Previous defenses, such as MetaCloak Liu et al. (2024) and SimAC Wang et al. (2024a), involve transformations like random cropping or Gaussian blur in the perturbation generation process to enhance robustness against preprocessing methods. However, these approaches fail to counter attacks like DiffPure Nie et al. (2022) and Noise Upscaling Hönig et al. (2025), which use diffusion models as noise purifiers. Directly incorporating Noise Upscaling or DiffPure into the optimization process can result in an excessively large computation graph. To address this challenge, we first apply small noises to $x_{p}$ according to the DiffPure and Noise Upscale settings and then maximize the denoising error loss across a set of substitute DiffPure and upscaling (super-resolution) models. The upscale loss function is defined as
@@ 4.4 The Algorithm of StyleGuard
A straightforward approach to tackle the bilevel problem in Section 4.1 is to unroll all training steps and optimize the protected examples $X_{p}$ through backpropagation. However, this would cause a very large computation graph that would exceed the capacity of most current machines. To overcome this challenge, inspired by Anti-DreamBooth Van Le et al. (2023), we use an approximate method to optimize the $X_{p}$ and $\theta$ alternatively. Specifically, in the $t$-th iteration, when the current model weights $\theta_{t}$ and the protected image set $X_{p}^{t}$ are available (with $\theta_{0}$ initialized from a pretrained diffusion model and $X_{p}^{0}:=X_{c}$), we create a copy of the current model weights, denoted as $\theta^{\prime}_{t,0}\leftarrow\theta_{t}$, for noise crafting. We then optimize the UNet of LDM for $K_{1}$ steps using DreamBooth loss:
@@
where $\eta$ and $\lambda$ are hyperpameters. We set $\eta$=1 and $\lambda$=10 in our experiments. Moreover, to improve the robustness to other transformations like crop and resize, we involve random transformations $\mathbb{T}^{\prime}$ in the optimization process based on the expectation over transformation (EoT) Liu et al. (2024); Athalye et al. (2018). Then we update the perturbed image $X_{p}$ using PGD with the attack budget $B_{\infty}$ for $K_{2}$ steps through
$X_{p}^{t+1}=\mathbb{E}_{g\sim\mathbb{T^{\prime}}}\left[\Pi_{B_{\infty}}\left(X_{p}^{t}+\alpha\,\text{sign}\left(\nabla_{X_{p}^{t}}L_{\text{styleguard}}\right)\right)\right],$ (8)
@@ 5.1 Experiment Setup
Initially, we use SD v1.4 and SD v1.5 as the substitute models to perturb the images. The attack budget is set as $\frac{8}{255}$ to be same with baselines. The images encoders used to compute the style loss includes VAE, OpenCLIP-ViT-H-14 and OpenCLIP-ViT-bigG-14. The StyleGuard training details and hyperparameters are included in Appendix A.1. During testing, we evaluate two popular fine-tuning methods: DreamBooth and Textual Inversion Gal et al. (2022b). For DreamBooth, we further assess two common settings: full-tuning (Full-FT) and LoRA fine-tuning (LoRA-FT) Hu et al. (2022), applied to both SD v2.1 and SD-XL (only LoRA on SD-XL, due to memory constraints). We used the official script from the Diffusers library for DreamBooth fine-tuning. The training details for DreamBooth and Textual Inversion are included in Appendix A.2 and A.3.
@@
We consider three kinds of attacks, including random transformation, DiffPure, and Noise Upscaling. For random transformations, we consider Gaussian noising, center cropping and resizing. For the DiffPure, we use the official code, and use Guided Diffusion Model and DDPM model during training, and use Stable Diffusion XL during evaluation. For Noise Upscaling, we followed the settings outlined by Honig et al. Hönig et al. (2025). During the optimization, we generate perturbation on the SD-x4-upscaler Rombach et al. (2022); stabilityai (2024b). During testing, we evaluate the defense effectiveness using the SD-x2-latent-upscaler stabilityai (2024a), which has a different architecture compared to the SD-x4-upscaler (see Appendix A.5 for more details). The number of diffusion steps was set to 30 because large diffusion steps will modify the image content. We also use an online black-box upscaler, Finegrain Image Enhancer (FIE) HuggingFace (2024). For FIE, the upscale factor is set as 2, and the ControlNet scale is set as 0.6.
@@
We compare our method with Glaze Shan et al. (2023), Mist Liang and Wu (2023), Anti-DreamBooth Van Le et al. (2023), MetaCloak Liu et al. (2024) and SimAC Wang et al. (2024a). The attack budget are set as $\frac{8}{255}$ for all baselines except Glaze, which uses LPIPS as constraint. The implementation details are included in Appendix A.4.
@@
To evaluate the defense performance for style mimicry, we use Fréchet Inception Distance (FID)Heusel et al. (2018) and precision Kynkäänniemi et al. (2019) as assessment metrics (consistent with MistLiang and Wu (2023)) and mimicry success rate Hönig et al. (2025), which uses human as annotators. To exclude content influence in the style mimicry attack, we use the model trained on clean images to generate 100 paintings in a specific category (e.g., an <sks> painting of a house). We then generate another 100 images from the model trained on style-guarded images using the same prompt and compute the FID and precision between the two sets. For personalization attacks, we use identity match score (IMS) Van Le et al. (2023) to access the semantic closeness between faces. The metric details are shown in the Appendix A.6.

### v2

@@ Abstract
Recently, text-to-image diffusion models have been widely used for style mimicry and personalized customization through methods such as DreamBooth and Textual Inversion. This has raised concerns about intellectual property protection and the generation of deceptive content. Recent studies, such as Glaze and Anti-DreamBooth, have proposed using adversarial noise to protect images from these attacks. However, recent purification-based methods, such as DiffPure and Noise Upscaling, have successfully attacked these latest defenses, showing the vulnerabilities of these methods. Moreover, present methods show limited transferability across models, making them less effective against unknown text-to-image models. To address these issues, we propose a novel anti-mimicry method, StyleGuard. We propose a novel style loss that optimizes the style-related features in the latent space to make it deviate from the original image, which improves model-agnostic transferability. Additionally, to enhance the perturbation’s ability to bypass diffusion-based purification, we designed a novel upscale loss that involves ensemble purifiers and upscalers during training. Extensive experiments on the WikiArt and CelebA datasets demonstrate that StyleGuard outperforms existing methods in robustness against various transformations and purifications, effectively countering style mimicry in various models. Moreover, StyleGuard is effective on different style mimicry methods, including DreamBooth and Textual Inversion. The code is available at https://github.com/PolyLiYJ/StyleGuard.
@@ 1 Introduction
Diffusion models have demonstrated remarkable effectiveness across various applications, such as image generation dhariwal2021diffusion ; ho2020denoising , image editing kawar2023imagic ; yang2023paint ; zhang2023sine , and text-to-image synthesis xie2023boxdiff ; wang2024compositional . The emergence of diffusion models has significantly transformed the art industry. These models allow users to create detailed artwork from simple text prompts, a task that once required extensive time and skill from professional artists. However, these technologies have also raised concerns about copyrights and ethics. For example, Dreambooth ruiz2023dreambooth allows anyone to fine-tune an SD model to imitate the artistic style of an artist with just a few paintings and generate high-quality artwork. This seriously damages the intellectual property rights of artists.
@@
To tackle the challenges associated with unauthorized image usage in text-to-image generation, recent perturbation-based approaches have emerged. These methods are designed to subtly modify user images, rendering them "unlearnable" for malicious applications and disrupting the functionality of targeted diffusion models. For example, Anti-Dreambooth van2023anti shows some effectiveness by alternately training diffusion models and executing PGD attacks, but it is fragile against simple data transformations. MetaCloak liu2024metacloak builds on Anti-Dreambooth by incorporating simple transformations into the attack pipeline. SimAC wang2024simac improves MetaCloak by selecting time steps with the highest gradients to improve the stability of the training.
@@
Although these methods have demonstrated effectiveness against style mimicry, recent purification-based methods pose significant challenges to these protections. For example, DiffPure nie2022diffusion and Noise Upscaling honig2025adversarial utilize LDM-based purifiers and upscalers to purify noise and successfully bypass the defenses of these methods, as illustrated in Figure 1. These novel attacks highlight the vulnerabilities of current anti-customization techniques. Moreover, in the real world, the finetuning model may differ from the model used in the training phase, making cross-model transferability an important issue. In light of these limitations, there is a need for a more efficient method to generate protective noise that is resilient to diffusion-based purifications and has strong cross-model transferability.
@@
To solve these problems, we propose StyleGuard, a more effective method based on style perturbations to protect artists from unauthorized text-to-image diffusion-based style mimicry. To generate transferable perturbations, inspired by style transfer huang2017arbitrary and adversarial attacks in feature space xu2021towards , we disrupt style-related features in latent space by adding subtle noise to the fine-tuning images that is nearly imperceptible to the human eye. Our method prevents text-to-image models from accurately extracting the style features of fine-tuning images, leading to false style correlations and hindering attackers from mimicking the original image’s style. Moreover, compared to Mist liang2023mist that uses $L_{2}$ distance in the feature space, which is prone to causing local disturbances and sensitive to image variations, style perturbations alter global features. As a result, our method is more robust to transformations such as image cropping, rescaling, and Gaussian noise. Additionally, bypassing diffusion-based purifications presents significant challenges, particularly because directly incorporating purifiers or upscalers into the optimization process can lead to memory overflow, as most purifiers and upscalers are based on diffusion models nie2022diffusion ; lee2023robust ; chen2024diffilter . To address this issue, we propose a novel upscale loss function that maximizes the loss of denoise error on ensemble purifiers and upscalers through a meta-learning approach. On the WikiArt and ClebA-HQ datasets, we show that StyleGuard significantly outperforms existing approaches over various transformations and purification methods and has higher cross-model transferability. Our main contributions are summarized as follows.
  • We propose StyleGuard, a robust method designed to effectively protect artists from DreamBooth-based style mimicry. StyleGuard accounts for various preprocessing techniques that attackers may employ, enhancing its practical effectiveness.
@@
  • To improve the model-agnostic transferability, we introduce a novel style loss, which aligns the style characteristics of the protected image more closely with those of the target image, allowing the fine-tuned model to establish an incorrect style connection.
@@
  • Experiments on the WikiArt and CelebA datasets demonstrate that StyleGuard offers enhanced protection against style mimicry and identity customization.
  • Compared to previous defense methods, our approach shows improved efficacy and practical effectiveness by considering various preprocessing techniques and model-agnostic scenarios.
@@ 2.1 Style Mimicry and Copyright Concerns
Unauthorized style mimicry has become a significant concern in the AI art community, where malicious actors exploit AI models to replicate an artist’s unique style without consent ANDERSEN2022 ; BAIO2022 ; MURPHY2022 . Such attacks often begin with a naive approach, where a generic text-to-image model is queried using the name of a well-known artist. More advanced mimicry attacks involve fine-tuning generic text-to-image models on a small collection of an artist’s works, often as few as 20 pieces by models such as DreamBooth gal2022image ; ruiz2023dreambooth . DreamBooth identifies key stylistic features and associates them with a specific token in the fine-tuned model, enabling highly accurate style replication. Such techniques have led to widespread incidents of unauthorized mimicry ANDERSEN2022 ; BAIO2022 ; MURPHY2022 , for example, CivitAI CIVITAI2022 built a large online website where people share their finetuned stable diffusion models. The potential for unauthorized style mimicry threatens the livelihoods of artists, leading to discussions about intellectual property rights in the digital age.
@@ 2.2 Protection Against Style Mimicry and Personalization
To address the unauthorized style mimicry issue, perturbation-based methods have been developed, which add subtle image perturbations to the unprotected images to disrupt generative models. For example, PhotoGuard salman2023raising aligns protected images’ latent features with black-and-white images. Glaze shan2023glaze minimizes the feature distance between perturbed images and target images while maintaining perceptual similarity. AdvDM liang2023adversarial reduces the likelihood of perturbed images under pre-trained diffusion models by disrupting the denoising process. Its enhanced version, Mist liang2023mist , utilizes black-and-white periodic images as targets and incorporates semantic loss and textual loss to improve protective strength. However, Mist directly minimizes the L2 distance between the original and target latent features, which causes noticeable and unnatural textures that degrade the original image’s quality. Anti-DreamBooth van2023anti proposes a novel scheme to defend the personalization attack that alternately updates the diffusion model and protected images. MetaCloak liu2024metacloak builds upon Anti-DreamBooth by incorporating simple transformations, such as Gaussian blur and cropping, into the attack pipeline to improve robustness against these transformations. SimAC wang2024simac further extends the work done by Anti-DreamBooth by selecting timesteps with maximum gradients to stabilize the training process. Some other methods huang2024disentangled ; zhu2024watermark ; yao2024promptcare ; wang2024invisible protect copyrights by embedding watermarks into images, subtly incorporating the author’s information. However, this kind of approach introduces additional verification processes.
While these methods exhibit robustness to simple transformations, recent work has introduced purification-based methods, including DiffPure nie2022diffusion and Noise Upscaling honig2025adversarial , which first add noise to images and then employ an LDM as purifier or upscaler to remove noise. These approaches have demonstrated impressive results in removing protective noise, rendering many recent protective methods ineffective. Therefore, there is an urgent need to develop more robust methods to defend against such attacks.
@@ 3.1 Style Mimicry by Dreambooth
DreamBooth ruiz2023dreambooth introduces a novel approach for personalizing text-to-image diffusion models by enabling them to generate high-fidelity images of specific subjects based on a few reference images. The method fine-tunes a pre-trained text-to-image model to bind a unique identifier (e.g., "[V]") to the subject, allowing the model to synthesize the subject in diverse contexts while preserving its key visual features. This is achieved through a fine-tuning process: the latent diffusion model (LDM) is fine-tuned using input images paired with text prompts containing the unique identifier and the subject’s class name (e.g., "An [V] painting"), while a class-specific prior preservation loss ensures the model retains its semantic understanding of the broader class (e.g., "A painting"). The DreamBooth loss function is defined as
@@ 4.2 Disturbing Style-related Features and Generating Transferable Perturbations
Previous work has found that the mean and variance of the feature space encapsulate the style information ulyanov2016instance ; xu2021towards ; huang2017arbitrarystyletransferrealtime ; karras2019stylebased . We consider the problem defined in Eq. 4.1 as maximizing the style-related feature distance between the reference images and the perturbed images and making it closer to the target images. Therefore, we propose a novel style loss function, which is defined as
@@ 4.3 Why Previous Defenses are Ineffective to Noise Upscaling and DiffPure
Previous defenses, such as MetaCloak liu2024metacloak and SimAC wang2024simac , involve transformations like random cropping or Gaussian blur in the perturbation generation process to enhance robustness against preprocessing methods. However, these approaches fail to counter attacks like DiffPure nie2022diffusion and Noise Upscaling honig2025adversarial , which use diffusion models as noise purifiers. Directly incorporating Noise Upscaling or DiffPure into the optimization process can result in an excessively large computation graph. To address this challenge, we first apply small noises to $x_{p}$ according to the DiffPure and Noise Upscale settings and then maximize the denoising error loss across a set of substitute DiffPure and upscaling (super-resolution) models. The upscale loss function is defined as
@@ 4.4 The Algorithm of StyleGuard
A straightforward approach to tackle the bilevel problem in Section 4.1 is to unroll all training steps and optimize the protected examples $X_{p}$ through backpropagation. However, this would cause a very large computation graph that would exceed the capacity of most current machines. To overcome this challenge, inspired by Anti-DreamBooth van2023anti , we use an approximate method to optimize the $X_{p}$ and $\theta$ alternatively. Specifically, in the $t$-th iteration, when the current model weights $\theta_{t}$ and the protected image set $X_{p}^{t}$ are available (with $\theta_{0}$ initialized from a pretrained diffusion model and $X_{p}^{0}:=X_{c}$), we create a copy of the current model weights, denoted as $\theta^{\prime}_{t,0}\leftarrow\theta_{t}$, for noise crafting. We then optimize the UNet of LDM for $K_{1}$ steps using DreamBooth loss:
@@
where $\eta$ and $\lambda$ are hyperpameters. We set $\eta$=1 and $\lambda$=10 in our experiments. Moreover, to improve the robustness to other transformations like crop and resize, we involve random transformations $\mathbb{T}^{\prime}$ in the optimization process based on the expectation over transformation (EoT) liu2024metacloak ; athalye2018synthesizing . Then we update the perturbed image $X_{p}$ using PGD with the attack budget $B_{\infty}$ for $K_{2}$ steps through
$X_{p}^{t+1}=\mathbb{E}_{g\sim\mathbb{T^{\prime}}}\left[\Pi_{B_{\infty}}\left(X_{p}^{t}-\alpha\,\text{sign}\left(\nabla_{X_{p}^{t}}L_{\text{styleguard}}\right)\right)\right],$ (8)
@@ 5.1 Experiment Setup
Initially, we use SD v1.4 and SD v1.5 as the substitute models to perturb the images. The attack budget is set as $\frac{8}{255}$ to be same with baselines. The images encoders used to compute the style loss includes VAE, OpenCLIP-ViT-H-14 and OpenCLIP-ViT-bigG-14. The StyleGuard training details and hyperparameters are included in Appendix A.1. During testing, we evaluate two popular fine-tuning methods: DreamBooth and Textual Inversion gal2022imageworthwordpersonalizing . For DreamBooth, we further assess two common settings: full-tuning (Full-FT) and LoRA fine-tuning (LoRA-FT) hu2022lora , applied to both SD v2.1 and SD-XL (only LoRA on SD-XL, due to memory constraints). We used the official script from the Diffusers library for DreamBooth fine-tuning. The training details for DreamBooth and Textual Inversion are included in Appendix A.2 and A.3.
@@
We consider three kinds of attacks, including random transformation, DiffPure, and Noise Upscaling. For random transformations, we consider Gaussian noising, center cropping and resizing. For the DiffPure, we use the official code, and use Guided Diffusion Model and DDPM model during training, and use Stable Diffusion XL during evaluation. For Noise Upscaling, we followed the settings outlined by Honig et al. honig2025adversarial . During the optimization, we generate perturbation on the SD-x4-upscaler Rombach_2022_CVPR ; x4upscalerHugging . During testing, we evaluate the defense effectiveness using the SD-x2-latent-upscaler sdx2latentupscalerHugging , which has a different architecture compared to the SD-x4-upscaler (see Appendix A.5 for more details). The number of diffusion steps was set to 30 because large diffusion steps will modify the image content. We also use an online black-box upscaler, Finegrain Image Enhancer (FIE) huggingfaceFinegrainImage . For FIE, the upscale factor is set as 2, and the ControlNet scale is set as 0.6.
@@
We compare our method with Glaze shan2023glaze , Mist liang2023mist , Anti-DreamBooth van2023anti , MetaCloak liu2024metacloak and SimAC wang2024simac . The attack budget are set as $\frac{8}{255}$ for all baselines except Glaze, which uses LPIPS as constraint. The implementation details are included in Appendix A.4.
@@
To evaluate the defense performance for style mimicry, we use Fréchet Inception Distance (FID)heusel2018ganstrainedtimescaleupdate and precision kynkaanniemi2019improved as assessment metrics (consistent with Mistliang2023mist ) and mimicry success rate honig2025adversarial , which uses human as annotators. To exclude content influence in the style mimicry attack, we use the model trained on clean images to generate 100 paintings in a specific category (e.g., an <sks> painting of a house). We then generate another 100 images from the model trained on style-guarded images using the same prompt and compute the FID and precision between the two sets. For personalization attacks, we use identity match score (IMS) van2023anti to access the semantic closeness between faces. The metric details are shown in the Appendix A.6.